OpenStudy (anonymous):

The area of a circle is 10π^2. What is the circle's diameter. Round to the nearest hundred of the centimeter. HELP!!!!!!

OpenStudy (anonymous):

Area of a Circle formula = πr^2 Given are of Circle = 10π^2 Equate both values, 10π^2 = πr^2 Cancel pi 10π = r^2 SquareRoot(10pi) = r Diameter = 2*SquareRoot(10pi) = 2*SquareRoot(10*3.14) = 2*SquareRoot(31.4) Diameter = 2*5.603 = 11.206 cm

OpenStudy (anonymous):

Thank you so much! :)